OpenStudy (anonymous):

Suppose that the functions p and q are defined as follows. p(x)=2X + 2 q(x)=-2x^2 find ( q . p)(2)= (p . q)(2)=

OpenStudy (anonymous):

Could someone help me with it please?

OpenStudy (anonymous):

The answer is (q.p)(2)= -48 (p.q)(2)=-48 right?

hartnn (hartnn):

yup, \(\large f.g\) is multiplication of those 2 functions

hartnn (hartnn):

and when you plug in x=2, you will get that answer.

hartnn (hartnn):

i meant \(\large p.q\)

OpenStudy (anonymous):

Yah but in the question there is a circle not a dot... is it the same?

OpenStudy (anonymous):

it's like that (q O p)(2)

hartnn (hartnn):

ohhh, this \(\large (p \:o \:q)\)(2) yes, it is diffferent!

OpenStudy (anonymous):

yup it's like that

hartnn (hartnn):

it is call composition of functions, read as q of p of 2 -->(q o p) (2)

hartnn (hartnn):

first find p(2) what do you get ?

hartnn (hartnn):

plug in x=2 in p(x)

OpenStudy (anonymous):

6

OpenStudy (anonymous):

q(x)=-8

hartnn (hartnn):

correct! so we have \((q o p)(2)= q (p (2))\) but p(2) = 6, so, \((q o p)(2)= q (p (2))=q(6)\) got this ?

OpenStudy (anonymous):

so in the first one it will be like this -8(6)?

OpenStudy (anonymous):

or just q(6)

hartnn (hartnn):

nopes, to get q(6) we put x=6 in q(x) what do u get ?

OpenStudy (anonymous):

don't know yet hhhh

hartnn (hartnn):

like you put x=2 in p(x) to get 2(2)+2 = 6 put x=6 in q(x) to get -2(6^2) =... ?

OpenStudy (anonymous):

ohhh now i see

OpenStudy (anonymous):

1 min i will solve it

OpenStudy (anonymous):

so the first one=-72 second=-14 right?

hartnn (hartnn):

oh, wow! you learn fast :D absolutely correct \(\huge \checkmark\)

OpenStudy (anonymous):

Thanks a lot Hartnn u r great (:

hartnn (hartnn):

lol, welcome ^_^